OpenStudy (elusive):

what do you need to do to x=sec(t) to get 1/x=cos(t)?

OpenStudy (photon336):

you know how secant and cosine are related?

OpenStudy (elusive):

sec =1/cos

jhonyy9 (jhonyy9):

yes exactly so than you know that sec(t) = 1/cos(t) so than result that cos(t) = 1/sec(t) given that x = sec(t) so from above wrote you know that 1/sec(t) = cos(t) x= sec(t) =>x = 1/cos(t) so 1/x = cos(t) hope this will help you

satellite73 (satellite73):

for example \[\sec(t)=2\iff \cos(t)=\frac{1}{2}\]

OpenStudy (elusive):

for instance.. how would you solve arcsec2 ?

satellite73 (satellite73):

you are solving \[\sec(x)=2\] so you would solve it just like the one i wrote above

satellite73 (satellite73):

i.e. you need to solve \[\cos(t)=\frac{1}{2}\]

OpenStudy (elusive):

pi/3?

satellite73 (satellite73):

yes

OpenStudy (elusive):

oh thank you! you can test me if you want

zepdrix (zepdrix):

arctan(-1)=?

OpenStudy (elusive):

@zepdrix ..is it 3pi/4?

OpenStudy (elusive):

I don't know :(

zepdrix (zepdrix):

You have the right idea... but these inverse functions bring a long a new set of problems. The range of our inverse tangent function is -pi/2 to pi/2. so with arctan, your resulting angle should always lie in the 1st or 4th quadrant.

zepdrix (zepdrix):

3pi/4 doesn't lie in either of those quadrants, ya? :| There is another angle which gives us -1 when we take the tangent of it.